OpenStudy (anonymous):

Analytic Trig, can someone help me with a series of questions for my homework

iYuko (iyuko):

For Trig I suggest using the qualified helpers option

OpenStudy (anonymous):

\[(\sin \theta -\cos \theta +1)/\sin \theta+\cos \theta -1) = \sin \theta +1/\cos \theta\]

iYuko (iyuko):

http://prntscr.com/6ubsfz Here!

iYuko (iyuko):

--------------------------------------------------------------^^^^^^^^^^^^^

OpenStudy (anonymous):

Where do you get the tan and sec from?